### 26 Jan 2024 v1

@@ 5 The evolution of $z=1$ unbarred galaxies
In this section, we will explore the evolution of the non-axisymmetry structures in the unbarred galaxies at $z=1$ to determine whether they develop a stable bar or not. For this, we will focus only on disc and lenticular descendants. As in the previous section, we break into two groups of unbarred galaxies: (1) the descendants of unbarred galaxies at $z=1$ and form a bar later in time, becoming barred galaxies at $z=0$; (2) the descendants of unbarred galaxies at $z=1$ were not able to develop a stable bar and remain unbarred at $z=0$.
@@
We find that $59$ unbarred galaxies, accounting for $29$ per cent of the parent sample (the $z=1$ disc galaxies), develop a stable bar structure at any point between $z=1$ and $z=0$. On the other hand, we find that 36 disc and lenticular galaxies ($18$ per cent of the parent sample) do not possess a bar structure at redshift $z=0$ and they were enabled to develop a stable bar structure after $z=1$.
@@
Figure 12: The evolution of unbarred disc galaxies divided into 2 subsamples: unbarred galaxies at $z=1$ which become barred galaxies at $z=0$ (left column), and those that were not able to develop a stable bar structure at any point up to $z=0$ (right column). Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution.
@@
Figure 13: The evolution of unbarred lenticular galaxies divided into 2 subsamples as in Fig.12: unbarred galaxies at $z=1$ which become barred galaxies at $z=0$ (left column), those that were not able to develop a stable bar structure at any point up to $z=0$ (right column). Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution.
@@
The evolution of the non-axisymmetric structures is illustrated in Fig. 12 over the disc descendants. The first column depicts the evolution of the non-axisymmetric structures in the descendant galaxies that form a bar after $z=1$ and keeps it until the present epoch ($z=0$). It can be seen that the formation of bar structures in these galaxies occurs shortly after the initial Gyrs. It should be noted that, on average, the bar is comparatively weaker and smaller compared to the subsample of barred galaxies that have managed to retain their bar structure (see Fig. 6), which is in line with them forming later in time. In terms of morphology, no significant evolutionary changes in the stellar mass fractions of the galaxy components are observed in this subsample.
@@
Similarly, the bar structures in lenticular descendants are formed just after $z=1$ as seen in Fig. 13. It should be noted that in the case of lenticular galaxies, the presence of a bar structure occurs during a stage where the progenitor galaxies possess a significant cold disc component and a relatively small bulge component before the morphological transformation takes place, and on average, these bars are stronger and longer than their disc counterparts.
@@
The evolution of the $A_{2,\rm max}$ can be seen in the second column of Fig. 12 for unbarred disc galaxies that never develop a stable bar. It is interesting that these galaxies have a relatively tiny bulge mass fraction and a correspondingly high disc mass fraction. Furthermore, there seems to be limited evolution observed in the thick disc and pseudobulge components. In the case of lenticular galaxies, there is an increase in the thick disc fraction and a decrease in the thin disc, whereas there is no or small evolution in the pseudobulge and bulge components. This is shown in the left panel of Fig. 13.
@@
To understand why some unbarred galaxies develop a bar whereas there are others that do not, we will investigate the role of the environment in the next section.
@@ 5.1 The role of environment in unbarred galaxies
Previous studies have proposed that mergers and tidal interactions can be another potential formation channel of the bars (e.g. Łokas et al. 2016; Martinez-Valpuesta et al. 2016; Łokas 2018; Izquierdo-Villalba et al. 2022). In this section, we will investigate whether the environment has a role in the formation of a bar or not for unbarred galaxies.
@@
Figure 14: The fraction of unbarred galaxies that experienced a major/minor merger since $z=1$ for the disc and lenticular descendants. Error bars represent the Poisson errors of each sample. The highest fraction of major/minor mergers corresponds to those galaxies which form a bar, whereas the lowest ratios are presented by unbarred galaxies which were not able to form a bar.
@@
Fig. 14 illustrates the fraction of disc (darker blue) and lenticular (fainter blue) galaxies that have experienced major/ minor mergers since $z=1$ split into galaxies that form a bar and galaxies that do not form a bar. The figure shows that the merger fractions with the lowest values (less than $0.25$) for disc descendant galaxies are those that have never formed a stable bar. The maximum fraction of galaxies experiencing major mergers is observed when disc galaxies form a bar, with a value of $0.37$. The fraction of galaxies that have experienced a minor merger exhibits a similar but more moderate trend across all subsamples. In particular, galaxies that are descendants of discs that never formed a bar demonstrate the lowest fractions. Conversely, disc galaxies that develop a bar, exhibit the highest fractions. It should be noted that the fraction of galaxies experiencing a minor merger is relatively smaller. Our findings indicate a possible correlation between the formation of a bar and the merger histories of galaxies. Indeed, previous studies, have proposed that mergers can be another potential formation channel of the bars (e.g. Łokas et al. 2016; Martinez-Valpuesta et al. 2016; Łokas 2018).
@@
Regarding lenticular galaxies, it has been observed in previous sections that they have more active merger histories than disc galaxies (see Fig.4), characterised by an elevated fraction of galaxies undergoing major and minor mergers. However, we observe a similar trend as that is typically observed in disc galaxies except lenticulars that never developed a stable bar. In this particular case, galaxies that have not developed bars have the highest merger fraction ($0.60$). This increase may account for the disruption of the disc, as depicted in Fig 13, during later stages of unbarred lenticular descendants galaxies that never developed a bar structure.
@@
Now, we will closely examine galaxies to see if these major mergers could trigger bar formation. For that, we calculate the temporal zone of the merger influence, as stated in Section 2.5. The quantification of this influence can be achieved through the parameter $\eta_{\rm dyn,bf}$, as defined in Eq. 1. With this parameter, the discrepancy between the time of bar formation and the time of the most recent major/minor merger or the upcoming major/minor merger is computed relative to the dynamical time of the halo at the moment of bar formation.
@@
Figure 15: Distributions of $\eta_{\rm dyn,bf}$ defined in Eq 1 for unbarred galaxies that formed their bar and undergo a major merger. This expression asses if the disc galaxy experienced a recent merger or not at the moment of bar dissolution, and values between $-2,2$ indicate the galaxy is still influenced by a recent merger. 30 (20) per cent of the disc (lenticular) unbarred galaxies that formed a bar, experienced a major merger in less than 2 dynamical times.
@@
Fig. 15 depicts the distribution of $\eta_{\rm dyn,bf}$ for descendant galaxies that underwent a delayed formation of their barred structure and were subjected to at least one significant merger event. Based on the depicted data, it is evident that a minority, less than $30$ per cent, of unbarred galaxies that undergo the formation of a bar are influenced by a massive merger event. Similarly, in the case of lenticular galaxies, approximately $20$ per cent of them are affected by a major merger during the period of bar formation. It should be noted that in the context of lenticular galaxies, the occurrence of recent mergers took place slightly after the formation of bars, as shown by the negative values of $\eta_{\rm dyn,bf}$. This suggests that a small fraction of bars was caused by a merger. In order to establish a causal relationship, we incorporate a random sample of disc galaxies that have undergone a massive merger, but $\eta_{\rm dyn,bf}$ calculated with an identical distribution of the bar formation times. Based on the plot, it can be observed that the random distribution exhibits a distribution with the peak centred at higher values of 2, which reinforces our suggestion that the connection between bar formation and mergers is just a small fraction. We also explore the distribution of $\eta_{\rm dyn,bf}$ for minor mergers and repeat the same exercise as before, where a similar distinction was observed between the random samples and the descendent galaxies that underwent bar formation when minor mergers were taken into account.
@@
We also investigate if tidal interactions with massive neighbours can contribute to the formation of bars. Fig. 16 shows the distance to the massive neighbours of these galaxies, and we found that there is no significant difference between unbarred galaxies that develop a bar and the ones that do not. Also, the satellite fraction as a function of lookback time increases similarly for both categories.
@@
Figure 16: The distance to a massive companion (mass ratio $=1/4$) of disc and lenticular descendant unbarred galaxies that form bar and do not undergo a major merger (orange line) versus unbarred galaxies that did not undergo a major merger, and they do not form a bar as a function of time since $z=1$. The environment seems to not play a role in the formation or not of a bar.
5.2 Why some unbarred galaxies do not develop a bar?
@@
This section focuses on analysing the galaxy properties associated with the formation of a bar. Specifically, we investigate the stellar mass, gas mass fraction, and the ratio of stellar mass to dark matter mass within a specific aperture (defined as two times the radius of the half stellar mass) for both disc and lenticular descendants. At the present time ($z=0$), galaxies with a stellar bar exhibit distinct characteristics compared to galaxies without bars. These differences include higher stellar mass (e.g Sheth et al. 2008) and lower gas fraction (e.g. Masters et al. 2012) and dominance of baryonic matter (e.g. Fragkoudi et al. 2021; Rosas-Guevara et al. 2022; Izquierdo-Villalba et al. 2022). The latter exhibits a notable disparity. We investigate this in more detail in Fig. 17.
@@
The diagram illustrates the relation between stellar mass and dark matter in unbarred discs at $z=1$ (shown in the bottom panels), as well as their descendants in the form of discs and lenticular galaxies at $z=0$ (top panels). The galaxies are divided into two groups: those that formed a stable bar (shown in the right column) and those that did not. According to the figure, unbarred galaxies that eventually form a bar tend to have a higher stellar mass compared to those that do not. Furthermore, in a general sense, unbarred galaxies have a lower ratio of stellar mass to dark matter compared to galaxies with a bar at $z=0$. This disparity is particularly pronounced among the less massive galaxies ($M_{*}<10^{10.5}\rm M_{\sun}$) in the sample. At $z=1$, it is shown that the ratio drops for both subsamples. However, there is no substantial difference between galaxies that produce a bar and those that do not. These results are consistent with the findings of Izquierdo-Villalba et al. (2022), who conducted a study on a sample of disc galaxies at $z=0$ in the TNG100 and TNG50 simulations, investigating the stellar mass-dark matter mass ratios. The study consistently found differences in galaxies that have a bar and do not have a bar. The authors also discovered differences in the stellar mass-dark matter ratio at different radii, even before bars formed. The aforementioned observation is also consistent with the recent findings reported by Fragkoudi et al. (2021) in their investigation of barred galaxies using the Auriga simulations. where has been established that the stellar component has been the main contributor to the overall rotation curve since $z=0.5$.
@@
Interestingly, Fig. 17 also shows that galaxies with a bar have a higher frequency of active merger histories compared to galaxies without a bar, thereby supporting our earlier findings (see 14). The colour scheme in Fig. 17 corresponds to the average count of major mergers. Furthermore, at $z=1$, it is shown that galaxies that will undergo bar formation after $z=1$ exhibit a higher count of major mergers at early epochs. This result supports the idea that barred galaxies have an early assembly of barred galaxies in comparison with unbarred galaxies, as shown in Rosas-Guevara et al. (2022, 2020). Given that, the variation in the assembly history of the barred galaxies also affects the structural parameters of their disc and bulge, which in turn influences the creation of a bar. This topic will be investigated in a future project, as it is not relevant to the current topic.
@@
It is noteworthy to examine the correlation between the presence of a bar and the stellar mass in Fig. 17. This pattern has been identified in multiple studies, most notably in Sheth et al. (2008) using a large sample of discs from the Cosmic Evolution Survey (COSMOS). Nevertheless, the authors have found that galaxies with a stellar mass of $10^{10}\rm M_{\sun}$ at $z=1$ have a bar fraction that is comparatively lower than that in the most massive galaxies at this redshift. Conversely, galaxies with a stellar mass of $10^{10}\rm M_{\sun}$ at $z=0$ exhibit a bar fraction comparable to that of the most massive galaxies at $z=0$. This may be in opposition to the results presented in Fig. 17, which depicts unbarred galaxies that not only undergo an increase in mass but also fail to form a bar, suggesting that the massive disc ability to effectively form a bar is diminished. As we have shown previously, this could be attributed to the excess of dark matter in the centre. However, this could also be attributed to the simulation galaxy formation model, which might be deficient in including crucial physical processes at small scales that trigger bar formation. Also, it is important to acknowledge that we are not conducting an apple-to-apple comparison in this study since we have to take into account the potential bias that the samples may possess. However, it is worth studying in more detail in future work.
@@
Figure 17: $M_{*}/M_{\rm DM}$ as a function of the stellar mass for central disc and lenticular galaxies at $z=1$ and $z=0$. The colour scheme corresponds to the mean number of major mergers for each bin since $z=2$. The top left panel represents the disc descendants that do not develop a bar, whereas the top right panel represents those that develop a bar. The bottom row corresponds to those galaxies at redshift $z=1$. The plots show that barred disc descendants of unbarred galaxies at $z=1$ have more mergers and higher stellar-to-dark matter ratios than those galaxies that never developed a bar even before the bar was formed for a given stellar mass.

### 07 Feb 2024 v2

@@ 5 The evolution of $z=1$ unbarred galaxies
In this section, we explore the evolution of the non-axisymmetry structures in the unbarred galaxies at $z=1$ to determine whether or not they develop a stable bar. For this, we focus on disc and lenticular descendants. As in the previous section, we break into two groups of unbarred galaxies: (1) the descendants of unbarred galaxies at $z=1$ that form a bar later in time, becoming barred galaxies at $z=0$; and (2) the descendants of unbarred galaxies at $z=1$ that were not able to develop a stable bar and remain unbarred at $z=0$.
@@
We find that $59$ unbarred galaxies, accounting for $29$ per cent of the parent sample (the $z=1$ disc galaxies), develop a stable bar structure at any point between $z=1$ and $z=0$. On the other hand, we find that 36 disc and lenticular galaxies ($18$ per cent of the parent sample) do not possess a bar structure at redshift $z=0$ but were able to develop a stable bar structure after $z=1$.
@@
Figure 12: Evolution of unbarred disc galaxies divided into two subsamples: unbarred galaxies at $z=1$ that become barred galaxies at $z=0$ (left column), and those that were not able to develop a stable bar structure at any point up to $z=0$ (right column). Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution.
@@
Figure 13: Evolution of unbarred lenticular galaxies divided into two subsamples as in Fig.12: unbarred galaxies at $z=1$ that became barred galaxies at $z=0$ (left column), and those that were not able to develop a stable bar structure at any point up to $z=0$ (right column). Solid lines represent the median values and the shaded area corresponds to the $20^{\rm th}$ and $80^{\rm th}$ percentiles of the distribution.
@@
The evolution of the non-axisymmetric structures is illustrated in Fig. 12 over the disc descendants. The first column depicts the evolution of the non-axisymmetric structures in the descendant galaxies that form a bar after $z=1$ and keep it until the present epoch ($z=0$). It can be seen that the formation of bar structures in these galaxies occurs shortly after the first few gigayears. It should be noted that, on average, the bar is comparatively weaker and smaller compared to the subsample of barred galaxies that have managed to retain their bar structure (see Fig. 6), which is in line with them forming later in time. In terms of morphology, no significant evolutionary changes in the stellar mass fractions of the galaxy components are observed in this subsample.
@@
Similarly, the bar structures in lenticular descendants are formed just after $z=1$ as seen in Fig. 13. It should be noted that in the case of lenticular galaxies, the presence of a bar structure occurs during a stage where the progenitor galaxies possess a significant cold-disc component and a relatively small bulge component before the morphological transformation takes place, and on average, these bars are stronger and longer than counterparts hosted by disc galaxies.
@@
The evolution of the $A_{2,\rm max}$ can be seen in the second column of Fig. 12 for unbarred disc galaxies that never develop a stable bar. It is interesting that these galaxies have a relatively tiny bulge mass fraction and a correspondingly high disc mass fraction. Furthermore, there seems to be limited evolution observed in the thick disc and pseudobulge components. In the case of lenticular galaxies, there is an increase in the thick-disc fraction and a decrease in the thin-disc fraction, whereas there is no or little evolution in the pseudobulge and bulge components. This is shown in the left panel of Fig. 13.
@@
To understand why some unbarred galaxies develop a bar whereas there are others that do not, we investigated the role of the environment. We present our findings in the following section.
@@ 5.1 The role of environment in unbarred galaxies
Previous studies proposed that mergers and tidal interactions can be another potential formation channel of bars (e.g. Łokas et al. 2016; Martinez-Valpuesta et al. 2016; Łokas 2018; Izquierdo-Villalba et al. 2022). In this section, we investigate whether or not environment has any influence on the likelihood that an unbarred galaxy will form a bar.
@@
Figure 14: Fraction of unbarred galaxies that experienced a major or minor merger since $z=1$ for the disc and lenticular descendants. Error bars represent the Poisson errors of each sample. The highest fraction of major or minor mergers corresponds to those galaxies that form a bar, whereas the lowest ratios are presented by unbarred galaxies that were not able to form a bar.
@@
Figure 14 illustrates the fraction of disc (darker blue) and lenticular (fainter blue) galaxies that have experienced major or minor mergers since $z=1,$ split into galaxies that form a bar and galaxies that do not form a bar. The figure shows that the lowest merger fractions (less than $0.25$) for disc descendant galaxies are shown by those that have never formed a stable bar. The maximum fraction of galaxies experiencing major mergers is observed when disc galaxies form a bar, with a value of $0.37$. The fraction of galaxies that have experienced a minor merger exhibit a similar but more moderate trend across all subsamples. In particular, galaxies that are descendants of discs that never formed a bar demonstrate the lowest fractions. Conversely, disc galaxies that develop a bar exhibit the highest fractions. It should be noted that the fraction of galaxies that experience a minor merger is relatively small. Our findings indicate a possible correlation between the merger history of a galaxy and the likelihood it will form a bar. Indeed, previous studies proposed that mergers can be another potential bar-formation channel (e.g. Łokas et al. 2016; Martinez-Valpuesta et al. 2016; Łokas 2018).
@@
As demonstrated in previous sections, lenticular galaxies have more active merger histories than disc galaxies (see Fig.4), characterised by an elevated fraction of galaxies undergoing major and minor mergers. However, we observe a similar trend to that typically observed in disc galaxies, except in lenticular galaxies that never developed a stable bar. In this particular case, galaxies that have not developed bars have the highest merger fraction ($0.60$). This increase may account for the disruption of the disc (as depicted in Fig. 13) during later stages of unbarred lenticular descendant galaxies that never developed a bar structure.
@@
Now, we closely examine galaxies to decipher whether or not these major mergers could trigger bar formation. To this end, we calculated the temporal zone of the merger influence, as stated in Section 2.5. The quantification of this influence can be achieved through the parameter $\eta_{\rm dyn,bf}$, as defined in Eq. 1. With this parameter, the discrepancy between the time of bar formation and the time of the most recent major or minor merger or the upcoming major or minor merger is computed relative to the dynamical time of the halo at the moment of bar formation.
@@
Figure 15: Distributions of $\eta_{\rm dyn,bf}$ defined in Eq. 1 for unbarred galaxies that formed their bar and underwent a major merger. This expression assesses whether or not the disc galaxy experienced a recent merger at the moment of bar dissolution, and values between $-2$ and $2$ indicate that the galaxy is still influenced by a recent merger. We find that 30 (20) per cent of the disc (lenticular) unbarred galaxies that formed a bar experienced a major merger within less than two dynamical times.
@@
Figure 15 depicts the distribution of $\eta_{\rm dyn,bf}$ for descendant galaxies that underwent delayed bar formation and were subjected to at least one significant merger event. Based on the depicted data, it is evident that the minority —that is, less than $30$ per cent— of unbarred galaxies that undergo bar formation are influenced by a massive merger event. Similarly, approximately $20$ per cent of lenticular galaxies are affected by a major merger during the period of bar formation. It should be noted that in the context of lenticular galaxies, the occurrence of a recent merger took place slightly after the formation of a bar, as shown by the negative values of $\eta_{\rm dyn,bf}$. This suggests that a small fraction of bars were caused by a merger. In order to establish a causal relationship, we incorporate a random sample of disc galaxies that have undergone a massive merger, but $\eta_{\rm dyn,bf}$ calculated with an identical distribution of the bar formation times. Based on the plot shown in Fig. 15, it can be observed that the random distribution exhibits a peak centred at values of higher than two, which reinforces our suggestion that a connection between bar formation and merger events is only seen in a small fraction of galaxies. We also explored the distribution of $\eta_{\rm dyn,bf}$ for minor mergers and repeated the same exercise as before, where a similar distinction was observed between the random samples and the descendent galaxies that underwent bar formation when minor mergers were taken into account.
@@
We also investigated whether or not tidal interactions with massive neighbours can contribute to the formation of bars. Figure 16 shows the distance to the massive neighbours of these galaxies, and we find that there is no significant difference between unbarred galaxies that develop a bar and those that do not. Also, the satellite fraction as a function of lookback time increases similarly for both categories.
@@
Figure 16: Distance to a massive companion (mass ratio $=1/4$) of disc and lenticular descendant unbarred galaxies that formed bar and did not undergo a major merger (orange line) versus unbarred galaxies that did not undergo a major merger as a function of time since $z=1$. The environment does not seem to not play a role in the likelihood that a disc or lenticular galaxy will form a bar.
5.2 Why do some unbarred galaxies not develop a bar?
@@
This section focuses on analysing the galaxy properties associated with the formation of a bar. Specifically, we investigated the stellar mass, gas mass fraction, and the ratio of stellar mass to dark matter mass within a specific aperture (defined as two times the radius of the half stellar mass) for both disc and lenticular descendants. At the present time ($z=0$), galaxies with a stellar bar exhibit characteristics that are distinct from those of galaxies without bars. These differences include higher stellar mass (e.g. Sheth et al. 2008), lower gas fraction (e.g. Masters et al. 2012), and dominance of baryonic matter (e.g. Fragkoudi et al. 2021; Rosas-Guevara et al. 2022; Izquierdo-Villalba et al. 2022). The latter disparity is the most significant. We investigate this in more detail in Fig. 17.
@@
The diagram illustrates the relation between stellar mass and dark matter in unbarred discs at $z=1$ (shown in the bottom panels), as well as their descendants in the form of discs and lenticular galaxies at $z=0$ (top panels). The galaxies are divided into two groups: those that formed a stable bar (shown in the right column) and those that did not. According to the figure, unbarred galaxies that eventually form a bar tend to have a higher stellar mass than those that do not. Furthermore, in a general sense, unbarred galaxies have a lower ratio of stellar mass to dark matter compared to galaxies with a bar at $z=0$. This disparity is particularly pronounced among the less massive galaxies ($M_{*}<10^{10.5}\rm M_{\sun}$) in the sample. At $z=1$, the ratio drops for both subsamples. However, there is no substantial difference between galaxies that produce a bar and those that do not. These results are consistent with the findings of Izquierdo-Villalba et al. (2022), who conducted a study on a sample of disc galaxies at $z=0$ in the TNG100 and TNG50 simulations, investigating the stellar mass–dark matter mass ratio. The study consistently found differences between galaxies with and without a bar. The authors also discovered differences in the stellar mass–dark matter ratio at different radii, even before bars formed. The aforementioned observation is also consistent with the recent findings reported by Fragkoudi et al. (2021) in their investigation of barred galaxies using the Auriga simulations; these authors established that the stellar component has been the main contributor to the overall rotation curve since $z=0.5$.
@@
Interestingly, Fig. 17 also shows that galaxies with a bar are more likely to have had an active merger history than galaxies without a bar, thereby supporting our earlier findings (see Fig. 14). The colour scheme in Fig. 17 corresponds to the average count of major mergers. Furthermore, at $z=1$, Fig. 17 shows that galaxies that will undergo bar formation after $z=1$ exhibit a higher count of major mergers at early epochs. This result supports the idea that barred galaxies have an earlier assembly in comparison with unbarred galaxies, as shown in Rosas-Guevara et al. (2022, 2020). Therefore, the variation in the assembly history of the barred galaxies also affects the structural parameters of their disc and bulge, which in turn influence the creation of a bar. This topic will be investigated in a future project, as it is not relevant to the present discussion.
@@
There is a noteworthy correlation between the presence of a bar and the stellar mass, as shown in Fig. 17. This pattern has been identified in multiple studies, most notably in Sheth et al. (2008) using a large sample of discs from the Cosmic Evolution Survey (COSMOS). Nevertheless, the authors found that galaxies with a stellar mass of $10^{10}\rm M_{\sun}$ at $z=1$ have a bar fraction that is comparatively lower than that in the most massive galaxies at this redshift. Conversely, galaxies with a stellar mass of $10^{10}\rm M_{\sun}$ at $z=0$ exhibit a bar fraction comparable to that of the most massive galaxies at $z=0$. This may be in opposition to the results presented in Fig. 17, which depicts unbarred galaxies that not only undergo an increase in mass but also fail to form a bar, suggesting that the massive disc ability to effectively form a bar is diminished. As we show in previous sections, this could be attributed to the excess of dark matter in the centre of such galaxies. However, this could also be attributed to the simulation galaxy formation model, which might be deficient in including crucial physical processes at small scales that trigger bar formation. Also, it is important to acknowledge that we are not make a perfect comparison in this study, as we have to take into account the potential bias that the samples may possess. However, it is worth studying this topic further in more detail in future work.
@@
Figure 17: $M_{*}/M_{\rm DM}$ as a function of stellar mass for central disc and lenticular galaxies at $z=1$ and $z=0$. The colour scheme corresponds to the mean number of major mergers for each bin since $z=2$. The top left panel represents the disc descendants that do not develop a bar, whereas the top right panel represents those that develop a bar. The bottom row corresponds to those galaxies at redshift $z=1$. Lenticular and disc galaxies that form a bar after $z=1$ experience more mergers and have higher stellar-to-dark matter ratios than those galaxies that never form a bar for a given stellar mass.